OpenStudy (anonymous):

Find an equation for the line: through (–4, –8) and horizontal. a. y = –8 b. x = –8 c. x = –4 d. y = –4

myininaya (myininaya):

horizontal lines are where the y values are constantly one number so y=that number What is your y in the point given on this line ?

OpenStudy (anonymous):

y = mx + b since the line is horizontal it has no slope (m) y = 0x+b y = x + b b is the y intercept; y is the y coordinate y = 0x + b -8 = x + b -8 = -4 + b -8 = -4 + -4 -8 = -8 d

OpenStudy (anonymous):

So would the answer be A Y= -8

myininaya (myininaya):

yea the y-coordinate of your point is -8 and your line is horizontal so it is y=-8